OpenStudy (anonymous):

Help Suppose H(x)=(6x-5)^2 Find the 2 functions f and g such that (f o g)(x)=H(x) Neither function can be the identity function. (There may be more than one correct answer.)

OpenStudy (campbell_st):

why not g(x) = 6x - 5 f(x) = (x)^2

OpenStudy (anonymous):

oh okay!

OpenStudy (anonymous):

I hope your right :D

OpenStudy (campbell_st):

well here is the method what is f(g(x)) = (g(x))^2 = (6x - 5)^2

OpenStudy (anonymous):

right

OpenStudy (campbell_st):

its means find a function g(x) substitute it into a function f(x) so that you get H(x) = (6x - 5)^2 you could have g(x) = (6x - 5)^2/3 f(x) = x^3 you'll get the same answer.

OpenStudy (anonymous):

but you said x^2 not x^3

OpenStudy (anonymous):

oh nevermind haha. that was an example right?

OpenStudy (campbell_st):

correct.... thats why the question says... there may be more than 1 solution.

OpenStudy (anonymous):

okay so the answer is g(x) =6x-5 and f(x) x^2... thanks